# Dealing with unbounded gradients in stochastic saddle-point optimization

### v1

@@ A.2.2 Eliminating the gradient norms
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\frac{\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)}{\eta_{y}T}+\frac{\eta_{y}}{2\gamma_{y}T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{y}(t)\right\|_{y,*}^{2}\right]+\frac{\mathcal{D}_{y}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)}{\eta_{x}T}+\frac{\eta_{x}}{2\gamma_{x}T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{x}(t)\right\|_{x,*}^{2}\right]$
@@
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\frac{\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)}{\eta_{y}T}+\frac{\eta_{y}}{2\gamma_{y}}L^{2}+\frac{\varrho_{y}}{2}\mathbb{E}\left[\left\|\bm{y}^{*}-\bm{y}_{1}\right\|^{2}_{x,*}\right]$
$\quad+\frac{\mathcal{D}_{y}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)}{\eta_{x}T}+\frac{\eta_{x}}{2\gamma_{x}}L^{2}+\frac{\varrho_{x}}{2}\mathbb{E}\left[\left\|\bm{x}^{*}-\bm{x}_{1}\right\|^{2}_{y,*}\right]$
$\quad+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\bm{y}_{t}-\bm{y}_{1}\right\|^{2}_{x,*}\right]\left(\frac{\eta_{x}L^{2}}{2\gamma_{x}}-\frac{\varrho_{y}}{2}\right)+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\bm{x}_{t}-\bm{x}_{1}\right\|^{2}_{y,*}\right]\left(\frac{\eta_{y}L^{2}}{2\gamma_{y}}-\frac{\varrho_{x}}{2}\right).$
Lastly, choosing $\varrho_{y}=\frac{\eta_{x}L^{2}}{\gamma_{x}}$ and $\varrho_{x}=\frac{\eta_{y}L^{2}}{\gamma_{y}}$ results in the bound stated in the theorem:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\frac{\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)}{\eta_{y}T}+\frac{\eta_{y}L^{2}}{2\gamma_{y}}+\frac{\varrho_{y}{\left\|\bm{y}^{*}-\bm{y}_{1}\right\|^{2}_{x,*}}}{2}$
$\quad+\frac{\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)}{\eta_{x}T}+\frac{\eta_{x}L^{2}}{2\gamma_{x}}+\frac{\varrho_{x}\left\|\bm{x}^{*}-\bm{x}_{1}\right\|^{2}_{y,*}}{2}.$
@@ B.2 Methodology
Let $\varrho_{v}=\eta_{\mu}$. Then, the output of COMIDA-MDP satisfies the following bound:
$\mathbb{E}\left[\left\langle\bm{\mu}^{\pi^{*}}-\bm{\mu}^{\overline{\pi}_{T}},\bm{r}\right\rangle\right]\leq\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{\pi^{*}}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}T}+\eta_{\mu}+\left(\frac{1}{\eta_{v}T}+4\eta_{\mu}\right)\mathbb{E}\left[\left\|{\bm{v}}^{\overline{\pi}_{T}}\right\|_{2}^{2}\right]+4\eta_{v}$
@@
By the standard online mirror descent analysis, we obtain the following upper bound on the first term that corresponds to the regret of the $\mu$-player:
@@
$\overset{(b)}{\leq}\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{*}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}}+\frac{\eta_{\mu}}{2}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{\mu}(t)\right\|_{\infty}^{2}\right]$ (20)
$\overset{(c)}{\leq}\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{*}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}}+\eta_{\mu}\sum_{t=1}^{T}\mathbb{E}\left[1+4\left\|{\bm{v}}_{t}\right\|_{\infty}^{2}\right]$
Here, we have used (a) Definition 2.2, (b) Corollary D.2 with $\mathcal{U}=\Delta_{SA}$, $\ell_{t}\left(\cdot\right)=-\mathcal{L}\left({\bm{v}}_{t};\cdot\right)$, $\mathcal{D}_{u}\!\left(\bm{u}\middle\|\bm{u}^{\prime}\right)=\mathcal{D}_{\textup{KL}}\left(\bm{u}\middle\|\bm{u}^{\prime}\right)$, $\varrho_{u}=0$ and $\bm{u}_{1}=\bm{\mu}_{1}$, as well as $(c)$ the bound on the gradient norm established in Equation (18).
As for the second term that corresponds to the regret of the $v$-player, the analysis is somewhat more involved. One challenge is that the comparator point ${\bm{v}}^{*}={\bm{v}}^{\overline{\pi}_{t}}$ is dependent on the iterates. This will be addressed at the end of this analysis. For now, applying Lemma D.1 with the appropriate parameters including $\mathcal{D}_{u}\!\left(\bm{u}\middle\|\bm{u}^{\prime}\right)=\frac{1}{2}\left\|\bm{u}-\bm{u}^{\prime}\right\|_{2}^{2}$, $\mathcal{H}_{u}\left(\bm{u}\right)=\mathcal{H}_{v}\left(\bm{u}\right)=\left\|\bm{u}\right\|_{\infty}^{2}$, $\bm{u}_{1}={\bm{v}}_{1}={\bm{{0}}}$ so that $\mathcal{H}_{u}\left(\bm{u}_{1}\right)=0$, as well as noting that the squared Euclidean norm is $1$-strongly convex and is the dual norm of itself gives the following bound:
@@
$\qquad\qquad\leq\frac{\mathbb{E}\left[\mathcal{D}_{v}\!\left({\bm{v}}^{*}\middle\|{\bm{v}}_{1}\right)\right]}{\eta_{v}}+\frac{\eta_{v}}{2}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{v}(t)\right\|_{2}^{2}\right]+\sum_{t=1}^{T}\mathbb{E}\left[\left\langle\bm{g}_{v}(t)-\widetilde{\bm{{g}}}_{v}(t),{\bm{v}}^{*}-{\bm{v}}_{t}\right\rangle\right]$
$\qquad\qquad\leq\frac{1}{\eta_{v}}\mathbb{E}\left[\mathcal{D}_{v}\!\left({\bm{v}}^{*}\middle\|{\bm{v}}_{1}\right)\right]+2\eta_{v}T+\sum_{t=1}^{T}\mathbb{E}\left[\left\langle\bm{g}_{v}(t)-\widetilde{\bm{{g}}}_{v}(t),{\bm{v}}^{*}-{\bm{v}}_{t}\right\rangle\right].$
@@
The last inequality follows from using that $\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{v}(t)\right\|_{2}^{2}\right]\leq 2$.
We are left with the problem of bounding the last term. We first observe that $\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\bm{g}_{v}(t)-\widetilde{\bm{{g}}}_{v}(t),{\bm{v}}_{t}\right\rangle\right]=0$ due to the $\mathcal{F}_{t-1}$-measurability of ${\bm{v}}_{t}$, so in fact all that remains is bounding $\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\bm{g}_{v}(t)-\widetilde{\bm{{g}}}_{v}(t),{\bm{v}}^{*}\right\rangle\right]$. This requires some care because ${\bm{v}}^{*}={\bm{v}}^{\overline{\pi}_{T}}$ depends on the entire sequence of iterates, and thus we cannot make direct use of the fact that $\mathbb{E}_{t}\left[\bm{g}_{v}(t)-\widetilde{\bm{{g}}}_{v}(t)\right]=0$. We will address this issue via a reduction to online learning inspired by the martingale tail bounds of Rakhlin & Sridharan (2017).
To this end, let us construct an auxiliary online learning game where in each round $t=1,2,\dots,T$, the following steps are repeated between an online learner and its environment:
  • The online learner chooses a function $\widehat{{\bm{v}}}_{t}\in^{S}$,
  • the environment chooses the cost function $\bm{c}_{t}=\widetilde{\bm{{g}}}_{v}(t)-\bm{g}_{v}(t)$,
  • the online learner incurs cost $\left\langle\bm{c}_{t},\widehat{{\bm{v}}}_{t}\right\rangle$ and observes $\bm{c}_{t}$.
By construction, we have that $\left\|\bm{c}_{t}\right\|_{2}\leq\left\|\bm{c}_{t}\right\|_{1}\leq 2$. We will study the case below where the online learner executes online gradient descent on the sequence $\left\{\bm{c}_{t}\right\}_{t=1}^{T}$, initialized at $\widehat{{\bm{v}}}_{1}={\bm{v}}_{1}$ and using stepsize $\eta_{v}$. By standard techniques (e.g., Lemma D.1 with $\varrho_{u}=0$ and), the regret of this method can be bounded against any comparator ${\bm{v}}^{*}$ as follows:
$\sum_{t=1}^{T}\left\langle\bm{c}_{t},\widehat{{\bm{v}}}_{t}-{\bm{v}}^{*}\right\rangle\leq\frac{\left\|{\bm{v}}^{*}-{\bm{v}}_{1}\right\|^{2}}{2\eta_{v}}+\frac{\eta_{v}}{2}\sum_{t=1}^{T}\left\|\bm{c}_{t}\right\|_{2}^{2}\leq\frac{\left\|{\bm{v}}^{*}-{\bm{v}}_{1}\right\|^{2}}{2\eta_{v}}+2\eta_{v}T.$
Then, the term we seek to bound can be controlled as follows:
$\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\bm{g}_{v}(t)-\widetilde{\bm{{g}}}_{v}(t),{\bm{v}}^{*}\right\rangle\right]$ $=-\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\bm{c}_{t},{\bm{v}}^{*}\right\rangle\right]=\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\bm{c}_{t},\widehat{{\bm{v}}}_{t}-{\bm{v}}^{*}\right\rangle\right]-\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\bm{c}_{t},\widehat{{\bm{v}}}_{t}\right\rangle\right]$
$\leq\frac{\mathbb{E}\left[\left\|{\bm{v}}^{*}-{\bm{v}}_{1}\right\|^{2}\right]}{2\eta_{v}}+2\eta_{v}T,$
where in the last step we used our regret bound stated just above, and also that $\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\bm{c}_{t},\widehat{{\bm{v}}}_{t}\right\rangle\right]=0$, which holds because the sum in question is a martingale. Indeed, note that for any $t$, we have
$\mathbb{E}_{t}\left[\left\langle\bm{c}_{t},\widehat{{\bm{v}}}_{t}\right\rangle\right]=\left\langle\mathbb{E}_{t}\left[\bm{c}_{t}\right],\widehat{{\bm{v}}}_{t}\right\rangle=\left\langle 0,\widehat{{\bm{v}}}_{t}\right\rangle=0,$
which holds due to the fact that $\widehat{{\bm{v}}}_{t}$ was chosen before $\bm{c}_{t}$ was revealed to the online learner. Overall, this proves
$\mathbb{E}\left[\sum_{t=1}^{T}\left\langle\bm{g}_{v}(t)-\widetilde{\bm{{g}}}_{v}(t),{\bm{v}}^{*}-{\bm{v}}_{t}\right\rangle\right]\leq\frac{\mathbb{E}\left[\left\|{\bm{v}}^{*}-{\bm{v}}_{1}\right\|^{2}\right]}{2\eta_{v}}+2\eta_{v}T,$
thus verifying the inequality
$\sum_{t=1}^{T}\mathbb{E}\left[\mathcal{L}^{\left(\text{reg}\right)}\left({\bm{v}}_{t};\bm{\mu}_{t}\right)-\mathcal{L}^{\left(\text{reg}\right)}\left({\bm{v}}^{*};\bm{\mu}_{t}\right)\right]\leq\frac{\mathbb{E}\left[\left\|{\bm{v}}^{*}-{\bm{v}}_{1}\right\|_{2}^{2}\right]}{\eta_{v}}+4\eta_{v}T.$
Putting the above inequality together with Equations (19) and (20), we finally obtain the following bound:
@@
$\mathbb{E}\left[G({\bm{v}}^{*};\bm{\mu}^{*})\right]$ $\leq\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{*}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}T}+\frac{\eta_{\mu}}{T}\sum_{t=1}^{T}\mathbb{E}\left[1+4\left\|{\bm{v}}_{t}\right\|_{\infty}^{2}\right]+\frac{\mathbb{E}\left[\left\|{\bm{v}}^{*}-{\bm{v}}_{1}\right\|_{2}^{2}\right]}{\eta_{v}T}+4\eta_{v}+\frac{\varrho_{v}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|{\bm{v}}^{*}\right\|_{\infty}^{2}-\left\|{\bm{v}}_{t}\right\|_{\infty}^{2}\right].$
@@ Appendix D Auxiliary Lemmas
[…]
$\qquad\qquad\leq\frac{\mathbb{E}\left[\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{1}\right)\right]}{\eta_{u}}+\frac{\eta_{u}}{2\gamma_{u}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{u}(t)\right\|_{u,*}^{2}\right]+\varrho_{u}\mathbb{E}\left[\mathcal{H}_{u}(\bm{u}_{1})\right]+\sum_{t=1}^{T}\mathbb{E}\left[\left\langle\bm{g}_{\bm{u}}(t)-\widetilde{\bm{{g}}}_{u}(t),\bm{u}^{*}-\bm{u}_{t}\right\rangle\right].$
@@
[…]
@@
$*.\displaystyle\qquad\qquad\leq\frac{1}{\eta_{u}}\Bigl{(}\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{t}\right)-\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{t+1}\right)\Bigr{)}+\frac{\eta_{u}}{2\gamma_{u}}\left\|\widetilde{\bm{{g}}}_{u}(t)\right\|_{u,*}^{2}+\left\langle\bm{g}_{\bm{u}}(t)-\widetilde{\bm{{g}}}_{u}(t),\bm{u}_{t}-\bm{u}^{*}\right\rangle.$
@@
$*\displaystyle\leq\frac{1}{\eta_{u}}\Bigl{(}\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{t}\right)-\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{t+1}\right)\Bigr{)}+\frac{\eta_{u}}{2\gamma_{u}}\left\|\widetilde{\bm{{g}}}_{u}(t)\right\|_{u,*}^{2}+\left\langle\bm{g}_{\bm{u}}(t)-\widetilde{\bm{{g}}}_{u}(t),\bm{u}_{t}-\bm{u}^{*}\right\rangle$
@@
Suppose the sequence of vectors $\{(\bm{x}_{t},\widetilde{\bm{{g}}}_{x}(t))\}_{t=1}^{T}$ is as described above. If the comparator $\bm{x}^{*}$ is fixed and independent of the iterates, the following inequality holds:
[…]
@@
Since $(\bm{u}_{t},\bm{y}_{t})$ is $\mathcal{F}_{t-1}$-measurable, $\mathbb{E}_{t}\left[\widetilde{\bm{{g}}}_{u}(t)\right]=\bm{g}_{\bm{u}}(t)$ and $\bm{u}^{*}$ does not depend on the iterates, the term $\left\langle\bm{g}_{\bm{u}}(t)-\widetilde{\bm{{g}}}_{u}(t),\bm{u}^{*}-\bm{u}_{t}\right\rangle$ is zero in expectation. Precisely,
@@
$\mathbb{E}\left[\left\langle\bm{g}_{\bm{u}}(t)-\widetilde{\bm{{g}}}_{u}(t),\bm{u}^{*}-\bm{u}_{t}\right\rangle\right]$ $=\mathbb{E}\left[\mathbb{E}_{t}\left[\left\langle\bm{g}_{\bm{u}}(t)-\widetilde{\bm{{g}}}_{u}(t),\bm{u}^{*}-\bm{u}_{t}\right\rangle\right]\right]$

### v2

@@ A.2.2 Eliminating the gradient norms
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\frac{2\mathbb{E}\left[\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)\right]}{\eta_{y}T}+\frac{\eta_{y}}{\gamma_{y}T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{y}(t)\right\|_{y,*}^{2}\right]+\frac{2\mathbb{E}\left[\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)\right]}{\eta_{x}T}+\frac{\eta_{x}}{\gamma_{x}T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{x}(t)\right\|_{x,*}^{2}\right]$
@@
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\frac{2\mathbb{E}\left[\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)\right]}{\eta_{y}T}+\frac{\eta_{y}}{\gamma_{y}}L^{2}+\frac{\varrho_{y}}{2}\mathbb{E}\left[\left\|\bm{y}^{*}-\bm{y}_{1}\right\|^{2}_{x,*}\right]$
$\quad+\frac{2\mathbb{E}\left[\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)\right]}{\eta_{x}T}+\frac{\eta_{x}}{\gamma_{x}}L^{2}+\frac{\varrho_{x}}{2}\mathbb{E}\left[\left\|\bm{x}^{*}-\bm{x}_{1}\right\|^{2}_{y,*}\right]$
$\quad+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\bm{y}_{t}-\bm{y}_{1}\right\|^{2}_{x,*}\right]\left(\frac{\eta_{x}L^{2}}{\gamma_{x}}-\frac{\varrho_{y}}{2}\right)+\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\bm{x}_{t}-\bm{x}_{1}\right\|^{2}_{y,*}\right]\left(\frac{\eta_{y}L^{2}}{\gamma_{y}}-\frac{\varrho_{x}}{2}\right).$
Lastly, choosing $\varrho_{y}=\frac{2\eta_{x}L^{2}}{\gamma_{x}}$ and $\varrho_{x}=\frac{2\eta_{y}L^{2}}{\gamma_{y}}$ results in the bound stated in the theorem:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\frac{2\mathbb{E}\left[\mathcal{D}_{y}\!\left(\bm{y}^{*}\middle\|\bm{y}_{1}\right)\right]}{\eta_{y}T}+\frac{\eta_{y}L^{2}}{\gamma_{y}}+\frac{\varrho_{y}}{2}\mathbb{E}\left[\left\|\bm{y}^{*}-\bm{y}_{1}\right\|^{2}_{x,*}\right]$
$\quad+\frac{2\mathbb{E}\left[\mathcal{D}_{x}\!\left(\bm{x}^{*}\middle\|\bm{x}_{1}\right)\right]}{\eta_{x}T}+\frac{\eta_{x}L^{2}}{\gamma_{x}}+\frac{\varrho_{x}}{2}\mathbb{E}\left[\left\|\bm{x}^{*}-\bm{x}_{1}\right\|^{2}_{y,*}\right].$
@@ B.2 Methodology
$\mathbb{E}\left[\left\langle\bm{\mu}^{\pi^{*}}-\bm{\mu}^{\overline{\pi}_{T}},\bm{r}\right\rangle\right]\leq\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{\pi^{*}}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}T}+\eta_{\mu}+\left(\frac{1}{\eta_{v}T}+4\eta_{\mu}\right)\mathbb{E}\left[\left\|{\bm{v}}^{\overline{\pi}_{T}}\right\|_{2}^{2}\right]+2\eta_{v}$
@@
Let $\bm{g}_{\mu}(t)=\nabla_{\bm{\mu}}\mathcal{L}({\bm{v}}_{t};\bm{\mu}_{t})$ denote the gradient of the Lagrangian in round $t$. By the standard online mirror descent analysis, we obtain the following upper bound on the first term that corresponds to the regret of the $\mu$-player:
@@
$\overset{(b)}{\leq}\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{*}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}}+\frac{\eta_{\mu}}{2}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{\mu}(t)\right\|_{\infty}^{2}\right]+\sum_{t=1}^{T}\mathbb{E}\left[\left\langle\bm{g}_{\mu}(t)-\widetilde{\bm{{g}}}_{\mu}(t),\bm{\mu}^{*}-\bm{\mu}_{t}\right\rangle\right]$
$\overset{(c)}{\leq}\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{*}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}}+\eta_{\mu}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{\mu}(t)\right\|_{\infty}^{2}\right]$ (20)
$\overset{(d)}{\leq}\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{*}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}}+\eta_{\mu}\sum_{t=1}^{T}\mathbb{E}\left[1+4\left\|{\bm{v}}_{t}\right\|_{\infty}^{2}\right]$
Here, we have used (a) Definition 2.2, (b) Lemma D.1 with $\mathcal{U}=\Delta_{SA}$, $\ell_{t}\left(\cdot\right)=-\mathcal{L}\left({\bm{v}}_{t};\cdot\right)$, $\mathcal{D}_{u}\!\left(\bm{u}\middle\|\bm{u}^{\prime}\right)=\mathcal{D}_{\textup{KL}}\left(\bm{u}\middle\|\bm{u}^{\prime}\right)$, $\varrho_{u}=0$ and $\bm{u}_{1}=\bm{\mu}_{1}$, $(c)$ that $\bm{\mu}^{*}$ is a fixed comparator and $\widetilde{\bm{{g}}}_{\mu}(t)$ is an unbiased estimate of $\bm{g}_{\mu}(t)$, as well as $(d)$ the bound on the gradient norm established in Equation (18).
As for the second term that corresponds to the regret of the $v$-player, the analysis is somewhat more involved. One challenge is that the comparator point ${\bm{v}}^{*}={\bm{v}}^{\overline{\pi}_{t}}$ is dependent on the iterates. To address this, we apply Corollary D.2 with the appropriate parameters including $\mathcal{U}=\mathbb{R}^{X}$, $\omega_{u}\left(u\right)=\frac{1}{2}\left\|\bm{u}\right\|_{2}^{2}$, , as well as noting that $\omega_{u}$ is $1$-strongly convex and is the dual norm of itself gives the following bound:
@@
$\qquad\qquad\leq\frac{\mathbb{E}\left[\left\|{\bm{v}}^{*}-{\bm{v}}_{1}\right\|_{2}^{2}\right]}{\eta_{v}}+\eta_{v}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{v}(t)\right\|_{2}^{2}\right]$ (21)
$\qquad\qquad\leq\frac{\mathbb{E}\left[\left\|{\bm{v}}^{*}-{\bm{v}}_{1}\right\|_{2}^{2}\right]}{\eta_{v}}+2\eta_{v}.$
@@
The last inequality follows from using that $\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{v}(t)\right\|_{2}^{2}\right]\leq 2$. Putting Equations (20) and (21) in Equation (19), we finally obtain the following bound:
$\rho^{*}-\rho^{\overline{\pi}_{T}}$ $=\mathbb{E}\left[G({\bm{v}}^{*};\bm{\mu}^{*})\right]$
@@
$\leq\frac{\mathcal{D}_{\textup{KL}}\left(\bm{\mu}^{*}\middle\|\bm{\mu}_{1}\right)}{\eta_{\mu}T}+\frac{\eta_{\mu}}{T}\sum_{t=1}^{T}\mathbb{E}\left[1+4\left\|{\bm{v}}_{t}\right\|_{\infty}^{2}\right]+\frac{\mathbb{E}\left[\left\|{\bm{v}}^{*}-{\bm{v}}_{1}\right\|_{2}^{2}\right]}{\eta_{v}T}+2\eta_{v}+\frac{\varrho_{v}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|{\bm{v}}^{*}\right\|_{\infty}^{2}-\left\|{\bm{v}}_{t}\right\|_{\infty}^{2}\right].$
@@ Appendix D Auxiliary Lemmas
$\qquad\qquad\leq\frac{\mathbb{E}\left[\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{1}\right)\right]}{\eta_{u}}+\frac{\eta_{u}}{2\gamma_{u}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{u}(t)\right\|_{u,*}^{2}\right]+\varrho_{u}\mathbb{E}\left[\mathcal{H}_{u}(\bm{u}_{1})\right]+\sum_{t=1}^{T}\mathbb{E}\left[\left\langle\bm{g}_{\bm{u}}(t)-\widetilde{\bm{{g}}}_{u}(t),\bm{u}_{t}-\bm{u}^{*}\right\rangle\right].$
@@
$\ell_{t}^{\left(\text{reg}\right)}\left(\bm{u}\right)=\ell_{t}\left(\bm{u}\right)+\varrho_{u}\mathcal{H}_{u}\left(\bm{u}\right).$ (23)
@@
$\qquad\qquad\leq\frac{1}{\eta_{u}}\Bigl{(}\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{t}\right)-\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{t+1}\right)\Bigr{)}+\frac{\eta_{u}}{2\gamma_{u}}\left\|\widetilde{\bm{{g}}}_{u}(t)\right\|_{u,*}^{2}+\left\langle\bm{g}_{\bm{u}}(t)-\widetilde{\bm{{g}}}_{u}(t),\bm{u}_{t}-\bm{u}^{*}\right\rangle.$
@@
$\leq\frac{1}{\eta_{u}}\Bigl{(}\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{t}\right)-\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{t+1}\right)\Bigr{)}+\frac{\eta_{u}}{2\gamma_{u}}\left\|\widetilde{\bm{{g}}}_{u}(t)\right\|_{u,*}^{2}+\left\langle\bm{g}_{\bm{u}}(t)-\widetilde{\bm{{g}}}_{u}(t),\bm{u}_{t}-\bm{u}^{*}\right\rangle$
@@
Suppose the sequence of vectors $\{(\bm{u}_{t},\widetilde{\bm{{g}}}_{u}(t))\}_{t=1}^{T}$ are as described in Lemma D.1 above. If the comparator $\bm{u}^{*}$ is adaptive and potentially dependent on the interaction history $\mathcal{F}_{T}$, the following inequality holds:
$\sum_{t=1}^{T}\mathbb{E}\left[\ell_{t}^{\left(\text{reg}\right)}\left(\bm{u}_{t}\right)-\ell_{t}^{\left(\text{reg}\right)}\left(\bm{u}^{*}\right)\right]\leq\frac{2\mathbb{E}\left[\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{1}\right)\right]}{\eta_{u}}+\frac{\eta_{u}}{\gamma_{u}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{u}(t)\right\|_{u,*}^{2}\right]+\varrho_{u}\mathbb{E}\left[\mathcal{H}_{u}(\bm{u}_{1})\right].$
@@
Recall that for any comparator $\bm{u}^{*}\in\mathcal{U}$ we can upper-bound the total regret from Lemma D.1 as:
@@
$\quad\leq\frac{\mathbb{E}\left[\mathcal{D}_{u}\!\left(\bm{u}^{*}\middle\|\bm{u}_{1}\right)\right]}{\eta_{u}}+\frac{\eta_{u}}{2\gamma_{u}}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{u}(t)\right\|_{u,*}^{2}\right]+\varrho_{u}\mathbb{E}\left[\mathcal{H}_{u}(\bm{u}_{1})\right]+\sum_{t=1}^{T}\mathbb{E}\left[\left\langle\bm{g}_{\bm{u}}(t)-\widetilde{\bm{{g}}}_{u}(t),\bm{u}_{t}-\bm{u}^{*}\right\rangle\right].$ (26)
Notice the last sum is zero when the comparator point $\bm{u}^{*}$ is independent of the interaction history $\mathcal{F}_{T}$, but this is generally not true for adaptively chosen comparators. A crude upper bound may be proved by treating it as a supremum of martingales, for instance by relying on the techniques of Rakhlin & Sridharan (2017). Inspired by their techniques, we provide a refined bound on this term based on a reduction to online learning, which has otherwise been first introduced in this context by Nemirovski et al. (2009). Lemma D.3 states the resulting bound which we highlight below: